OpenStudy (anonymous):

Explain what tranformations you would need to apply to the graph of y=f(x). y=-f(1/2x)-2

OpenStudy (sirm3d):

@lovekblue are you done reviewing?

OpenStudy (anonymous):

my book says it will be a horitonzal stretch by by a factor of two, but why? I do not understand

OpenStudy (anonymous):

-f(x) = flip on the x axis 1/2 x = horizontal stretch by 2 -2 = move down 2 units. is that right. .

OpenStudy (anonymous):

review what? my own exam stuff? totally not

OpenStudy (anonymous):

yes it is, but I dont understand why it would be a stretch by a factor of 2

OpenStudy (sirm3d):

you're right in your explanation, lovekblue

OpenStudy (anonymous):

oh yea, u explain it :D i gotta go back to my question page lol

OpenStudy (anonymous):

help me please?

OpenStudy (sirm3d):

if you use x=1 in the original function, the transform \[\frac{ 1 }{ 2 }x\] will use only 1/2

OpenStudy (sirm3d):

one unit length in the original function is equivalent to half unit in the transform, so the graph is stretched horizontally

OpenStudy (anonymous):

but why would the answer be 2 and not 1/2 ?

OpenStudy (anonymous):

or using my dumb way to explain... since x is being halved, u need to have a greater x input to get the y . so the graph stretches outward.

OpenStudy (sirm3d):

if on your graphing paper, the farthest point on the horizontal axis is 6, on the transform, that is only at 3

OpenStudy (anonymous):

let's say ur original function is x=1 , but now u have 1/2 x = 1 , to accomodate tha change, you need to multiply that by 2 to get y=1 isnt it? that's why you are changing that x by a factor of 2

OpenStudy (anonymous):

sirm, is my logic correct lol

OpenStudy (sirm3d):

if you try to make the two points (6,0) on the original axis and the point (3,0) on the transform, would you not stretch the transform axis just to make them coincide?

OpenStudy (sirm3d):

that is the restorative effect

OpenStudy (sirm3d):

which is true of course

OpenStudy (sirm3d):

which is true of course

OpenStudy (anonymous):

so when I multiply by 2 it would look like this? 2 x 1/2x= 1 ? then 2/2x=1 then wouldn't it be 1x=1?

OpenStudy (sirm3d):

that would be fine

OpenStudy (sirm3d):

if a transform is \[\frac{ 3 }{ 2 }x \], the restorative factor is \[\frac{ 2 }{ 3 } \] so the graph "shrinks" horizontally

OpenStudy (anonymous):

hm so suppose my graph said y=-f(2/3x)-2 would I say it would be a horitonzal compression by 2/3 or would it be something else?

OpenStudy (anonymous):

because for the last equation we did not say it was a stretch by 1/2, we said 2

OpenStudy (sirm3d):

horizontal compression it is

OpenStudy (sirm3d):

oops, still expansion

OpenStudy (sirm3d):

what factor must you multiply \[\frac{ 2 }{ 3 }x \] to get x? that factor will tell you its its compression (less than 1) or expansion (greater than 1)

OpenStudy (anonymous):

3/2 ?

OpenStudy (sirm3d):

yes, 3/2. and that is greater than 1, so a horizontal expansion

OpenStudy (anonymous):

Wow thank you sooo much! You helped me alot :) you and your friend. Thank you once again !

OpenStudy (sirm3d):

you seem to get lovekblue's idea more easily.

OpenStudy (anonymous):

lol cuz my brain is structured to process only easy things :D

OpenStudy (anonymous):

No you were both a big help :) aha god bless

OpenStudy (sirm3d):

easy things are easier to remember. haahaha